# CostFilter-AD: Enhancing Anomaly Detection through Matching Cost Filtering

### v1

@@ 4.4 Further Analysis
Time and memory efficiency. Table 6 compares the inference speed and memory usage per image on an A100 GPU (40GB), with and without our method. The results highlight the computational efficiency and minimal overhead of our plug-in design. Seamlessly integrating into reconstruction- or embedding-based pipelines, our method enables efficient and precise anomaly detection and localization, underscoring its practicality for real-world applications.
Table 6: Comparison of inference time and memory usage.
Method Time (s) Memory (GB)
GLAD / +Ours 3.96 / +0.37 8.79 / +2.07
HVQ-Trans / +Ours 0.05 / +0.07 4.78 / +0.94
AnomalyDF / +Ours 0.31 / +0.32 3.25 / +0.82
@@ Acknowledgements
This work is supported in part by the Research Program of the Liaoning Liaohe Laboratory under Grant No. LLL23ZZ-05-01, in part by the Natural Science Foundation of Liaoning Province of China under Grant No. 2024-MSBA-42, in part by the Key Research and Development Program of Liaoning Province under Grant No. 2023JH26/10200011, and in part by the Program of China Scholarship Council 202306080142.
@@
This paper presents work whose goal is to advance unsupervised anomaly detection through CostFilter-AD, a generic post-processing plug-in designed to address matching noise, which is widely present in existing methods. We reformulate anomaly detection as a three-step pipeline: feature extraction, cost volume construction, and cost volume filtering, providing a scalable and precise solution. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@
  • Sec. D reports comprehensive per-category quantitative results on the MVTec-AD dataset.
  • Sec. E reports comprehensive per-category quantitative results on the VisA dataset.
  • Sec. F showcases comprehensive qualitative results for each category on the MVTec-AD dataset.
  • Sec. G showcases comprehensive qualitative results for each category on the VisA dataset.
@@
(i) AnomalDF/AnomalDF+Ours. In this paper, we refer to the full-shot version of AnomalDINO (Damm et al., 2025) as AnomalDF, where “F” denotes the full-shot setting. AnomalDINO explores two experimental settings: few-shot and full-shot. The few-shot setting assumes access to only a limited number of normal templates per category, whereas the full-shot setting constructs a comprehensive memory bank of normal template features from the training set, following the experimental setup of PatchCore (Roth et al., 2022). However, this full-shot approach incurs substantial storage overhead.
To address this, we optimize global feature matching and denoising by leveraging a limited selection of normal templates. Specifically, each test image is matched globally against $N=3$ randomly selected normal templates, rather than relying on a fixed set as in the few-shot setting, defining AnomalDF. During training, our matching cost filtering model follows the same procedure: performing global feature matching between an input image and $N=3$ randomly selected normal templates, followed by anomaly cost volume construction and cost volume filtering. Through iterative training, the model progressively learns feature representations from all normal templates over multiple epochs, forming our enhanced approach, AnomalDF+Ours.
@@
(iii) Residual Channel-Spatial Attention (RCSA) module. Fig. 4 illustrates the RCSA module (Eq. 3.3), which generates two attention tensors: a channel attention tensor of shape $(B,C^{\prime},1,1,1)$ and a spatial attention tensor of shape $(B,1,D^{\prime},H^{\prime},W^{\prime})$, where $D^{\prime}$ represents the depth of features at each layer. These tensors refine feature representations across both matching and spatial dimensions. The RCSA module enhances global feature matching via residual channel attention and improves pixel-level anomaly localization through residual spatial attention. Notably, the residual connections preserve anomaly-relevant features, facilitating progressive denoising and precise anomaly detection.
(iv) Evaluation protocol. We adopt standard practice for qualitative visualization by plotting the abnormal map logits of a single image to ensure clearer comparisons. For quantitative evaluation and KDE logit curve generation, we normalize the normal-abnormal map logits across all images within a category, improving detection robustness and comparability.
@@ Appendix B Challenges in GLAD: Visual Analysis and Comparison with Our Method
(i) Asymmetry in multi-class reconstruction induces significant matching noise artifacts. As shown in Fig. 5 (a), a unified multi-class model must reconstruct diverse anomalies, frequently distorting shape, texture, or orientation, resulting in asymmetric feature matching and significant matching noise artifacts. This issue is also prevalent in embedding-based methods (Roth et al., 2022; Damm et al., 2025). Our approach mitigates it by leveraging input feature guidance within dual-stream attention, guiding the filtering model to focus on spatial structures while preserving edge details. This effectively enhances anomaly localization, making it compatible with both reconstruction-based and embedding-based models.

### v4

@@ 4.4 Further Analysis
Time and memory efficiency. Table 7 presents a comparative analysis of computational efficiency across multiple baselines, reporting parameter count, FLOPs, memory consumption, and per-image inference latency. All measurements were obtained with a batch size of 1 on an A100 GPU (40 GB), both with and without our method. The results demonstrate that our plug-in consistently improves performance while introducing only reasonable memory overhead and modest computational cost. Notably, the increase in memory usage is minimal, and the additional inference latency remains low, particularly when compared to diffusion-based approaches. Minor variations in #Params stem from projecting matching features with different channel dimensions (e.g., 196, 768, 1024) into a unified 96-dimensional space. By maintaining architectural compatibility and low integration overhead, our method enables efficient and accurate anomaly detection and localization, reinforcing its applicability to real-world deployment scenarios.
Table 7: Comprehensive comparison of baselines and + Ours in terms of parameter size (#Params), computational cost (FLOPs), memory usage (Mem.), and per-image inference time (Inf.).
Method #Params FLOPs Mem. (GB) Inf. (s/image)
UniAD / +Ours 7.7M / +43.0M 198.0G / +26.0G 4.53 / +0.56 0.01 / +0.04
Glad/+Ours 1.3B / +43.8M >>2.2T / +32.7G 8.79 / +2.07 3.96 / +0.37
HVQ-Trans/+Ours 18.0M / +43.0M 7.4G / +26.0G 4.78 / +0.94 0.05 / +0.07
AnomalDF/+Ours 21.0M / +43.8M 4.9G / +32.7G 3.25 / +0.82 0.31 / +0.32
Dinomaly/+Ours 132.8M / +43.6M 104.7G / +14.3G 4.32 / +1.11 0.11 / +0.05
4.5 Failure Case Analysis
As illustrated in Fig. 4, we present representative failure cases from six categories in MVTec-AD and VisA, demonstrating the results of our method before and after filtering. While our method effectively suppresses the matching noise, it mainly relies on the presence of anomaly-relevant signals in the cost volume. To address situations where such signals are limited due to low-resolution inputs or insufficient feature extraction, we introduce input image features for spatial guidance. Nevertheless, recovery remains challenging in these cases, indicating a promising direction for future work.
Figure 4: Failure cases on MVTec-AD and VisA. Some subtle anomalies may lead to inaccurate or missed localization, potentially due to limited representation in the constructed cost volume.
@@ Acknowledgements
This work is supported in part by the Research Program of the Liaoning Liaohe Laboratory under Grant No. LLL23ZZ-05-01, in part by the Natural Science Foundation of Liaoning Province of China under Grant No. 2024-MSBA-42, in part by the Key Research and Development Program of Liaoning Province under Grant No. 2023JH26/10200011, in part by the Science and Technology Major Project of Liaoning Province under Grant No. 2024JH1/11700048, in part by the Fundamental Research Funds for the Central Universities under Grant No. N25YJS002, in part by the Major Program of the National Natural Science Foundation of China (NSFC) under Grant No. 61991404, and in part by the Program of China Scholarship Council 202306080142.
@@
This paper presents work whose goal is to advance unsupervised anomaly detection through CostFilter-AD, a generic post-processing plug-in designed to address matching noise, which is widely present in existing methods. We reformulate anomaly detection as a three-step pipeline: feature extraction, cost volume construction, and cost volume filtering, providing a scalable and precise solution. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@
  • Sec. D provides evaluations of CostFilter-AD under multiple original AnomalyDINO full-shot settings.
  • Sec. E showcases comprehensive qualitative results for each category on the MVTec-AD dataset.
  • Sec. F showcases comprehensive qualitative results for each category on the VisA dataset.
  • Sec. G reports comprehensive per-category quantitative results on the MVTec-AD dataset.
  • Sec. H reports comprehensive per-category quantitative results on the VisA dataset.
  • Sec. I reports comprehensive per-category quantitative results on the BTAD dataset.
  • Sec. J reports comprehensive per-category quantitative results on the MPDD dataset.
@@
(i) AnomalDF/AnomalDF+Ours. In this paper, we refer to the full-shot version of AnomalDINO (Damm et al., 2025) as AnomalDF, where “F” denotes the full-shot setting. AnomalDINO explores two experimental settings: few-shot and full-shot. The few-shot setting assumes access to only a few fixed normal templates per category, whereas the full-shot setting constructs a big memory bank of all normal template features from the training set. However, this full-shot approach incurs more storage overhead.
To address this, we optimize global feature matching and denoising by leveraging a limited selection of normal templates. During training, AnomalDF (+Ours) uses a fixed number ($N=3$) of templates per input image, randomly sampled from the full training set for each input, rather than drawn from a fixed template set as in the original few-shot setting of AnomalyDINO. Our dynamic sampling ensures the template pool covers the full training distribution; thus, we classify it as full-shot, offering a trade-off between template diversity and memory efficiency. During testing, for fairness, we evaluate AnomalDF (+Ours) using our dynamic 3-shot sampling protocol, as reflected in the results reported in our submission.
@@
(iii) Evaluation protocol. We adopt standard practice for qualitative visualization by plotting the abnormal map logits of a single image to ensure clearer comparisons. For quantitative evaluation and KDE logit curve generation, we normalize the normal-abnormal map logits across all images within a category, improving detection robustness and comparability.
(iv) Utilization of input/reconstructed features in reconstruction-based methods. For reconstruction-based UAD methods such as HVQ-Trans, UniAD, and Dinomaly, we construct the cost volume directly using both the input and reconstructed features, without decoding them back to the image domain. This design is based on the fact that these methods already produce semantically meaningful representations in latent space. In other words, we leverage the multi-layer decoder features directly for cost volume construction, instead of relying on an external pre-trained encoder (e.g., DINO) as done in GLAD+Ours and AnomalDF+Ours.
(v) Residual Channel-Spatial Attention (RCSA) module. Fig. 5 illustrates the RCSA module (Eq. 3.3), which generates two attention tensors: a channel attention tensor of shape $(B,C^{\prime},1,1,1)$ and a spatial attention tensor of shape $(B,1,D^{\prime},H^{\prime},W^{\prime})$, where $D^{\prime}$ represents the depth of features at each layer. These tensors refine feature representations across both matching and spatial dimensions. The RCSA module enhances global feature matching via residual channel attention and improves pixel-level anomaly localization through residual spatial attention. Notably, the residual connections preserve anomaly-relevant features, facilitating progressive denoising and precise anomaly detection.
@@ Appendix B Challenges in GLAD: Visual Analysis and Comparison with Our Method
(i) Asymmetry in multi-class reconstruction induces significant matching noise artifacts. As shown in Fig. 6 (a), a unified multi-class model must reconstruct diverse anomalies, frequently distorting shape, texture, or orientation, resulting in asymmetric feature matching and significant matching noise artifacts. This issue is also prevalent in embedding-based methods (Roth et al., 2022; Damm et al., 2025; Zhong et al., 2022). Our approach mitigates it by leveraging input feature guidance within dual-stream attention, guiding the filtering model to focus on spatial structures while preserving edge details. This effectively enhances anomaly localization, making it compatible with both reconstruction-based and embedding-based models.
@@ Appendix C Progressive Matching Noise Suppression via CostFilter-AD
Table 8: AnomalDF vs. AnomalDF+Ours under a comprehensive setting on MVTec-AD and VisA.
ID Dataset Method Resize Temples I-AUROC I-AP I-F1max P-AUROC P-AP P-F1max P-AUPRO
1 MVTec-AD AnomalDF 256 3 96.8 98.6 97.1 98.1 61.3 60.8 93.6
2 +Ours 256 3 98.5 99.4 97.8 98.8 67.8 64.9 94.2
3 AnomalDF 256 Full 99.0 99.3 98.4 97.5 58.7 91.7
4 +Ours 256 Full 99.3 99.8 98.6 98.9 68.7 65.5 96.6
5 AnomalDF 448 Full 99.3 99.7 98.8 97.9 61.8 92.9
6 +Ours 448 Full 99.5 99.8 98.9 99.0 72.4 68.4 95.4
7 AnomalDF 672 Full 99.5 99.8 99.0 98.2 64.3 95.0
8 +Ours 672 Full 99.6 99.9 99.0 99.1 74.4 69.7 96.3
9 VisA AnomalDF 256 3 90.5 91.4 86.2 97.4 39.6 40.4 86.3
10 +Ours 256 3 94.3 95.1 90.6 99.2 44.6 45.5 84.5
11 AnomalDF 256 Full 94.6 95.7 90.9 98.3 44.3 86.7
12 +Ours 256 Full 95.5 96.3 91.5 99.4 45.9 46.6 87.0
13 AnomalDF 448 Full 97.2 97.6 93.7 98.7 50.5 95.0
14 +Ours 448 Full 97.4 97.7 93.8 99.4 42.2 53.6 95.2
15 AnomalDF 672 Full 97.6 97.2 94.3 98.9 53.8 96.1
16 +Ours 672 Full 97.8 98.0 94.6 99.4 47.6 54.5 96.4
Appendix D Evaluation Under the Original AnomalyDINO Full-shot Settings
In our main paper, AnomalDF and AnomalDF+Ours were trained with 3 randomly sampled reference templates per input and a resolution of $256\times 256$, and evaluated using a similarly limited number of templates, offering a trade-off between template diversity and memory efficiency. Exp. ID 1, 2, 9, and 10 in Table 8 report the corresponding results.
In contrast, the original full-shot setting of AnomalyDINO (Damm et al., 2025) utilizes the entire training set as reference templates and resizes images to larger resolution. To ensure a fair and thorough comparison, we further conducted evaluations under the original full-shot setting of AnomalyDINO. Exp. ID 3 to 8 and Exp. ID 11 to 16 in Table 8 report results on MVTec-AD and VisA, respectively. Notably, to mitigate storage and compute overhead, we directly reuse the models trained in Exp. ID 2 and 10, and test them under different resolutions and template amounts, without additional retraining. Despite this constraint, CostFilter-AD consistently improved the performance of AnomalyDINO across various resolutions and datasets. Remarkably, our method at lower resolution (e.g., $448\times 448$) may match or outperform the original AnomalDF baseline at higher resolution (e.g., $672\times 672$), demonstrating its effectiveness. These results highlight the scalability and generalization capability of our plug-in method, even under varied operational constraints.